OpenStudy (anonymous):

The sum of two consecutive even integers is 234. What is the larger integer? 114 116 118 124

hero (hero):

x + x + 2 = 234 2x + 2 = 234 2x = 232 x = 116 The larger is 118

hero (hero):

116 + 118 = 234